OpenStudy (anonymous):

Can anyone help me with algebra 2 please???? Im begging you?!?!?!?

OpenStudy (anonymous):

this is the question im so lost

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My birthday 12/04/1998

OpenStudy (vera_ewing):

f(x)=3(x-8)2+15

OpenStudy (anonymous):

is that for the first question??

OpenStudy (vera_ewing):

yes

OpenStudy (vera_ewing):

A. g(x)=2(x+-8)2+15. This meets the requirements because it opens in the same direction as function f(x) because a is positive which means it opens upward. It has the same vertex, 10, 1 because the h and k are the same as the function f(x), and it will be narrower since a is 2 which is less than the 3 used in the function f(x). B. h(x)=-2(x-8)2+15. This meets the requirements because it will open in the opposite direction, downwards, since a is negative. It has the same vertex, 10, 1 because the h and k are the same as the function f(x), and it will be wider because -2 is less than 3 and the smaller the number, the wider the parabola.

OpenStudy (vera_ewing):

Google how to change an equation from vertex to standard form for #2

OpenStudy (anonymous):

but doesnt it have to go by my birthday

OpenStudy (vera_ewing):

Yeah I'll give you a link that should help you :)

OpenStudy (anonymous):

yeah i looked at that but i have a diff bday so its not helpful

OpenStudy (vera_ewing):

do you actually have to use your birthday? does your teacher know your exact birthday? i'm sure she/he won't mind because your teacher just wants to make sure you now how to do it :)

OpenStudy (vera_ewing):

also, they show the steps and everything. just plug in your birthday if you want. :)

OpenStudy (anonymous):

yeah but im taking it on an online thing and itll know if i use someone else work. Plus i want to understand how to do it

OpenStudy (vera_ewing):

what online thing? :)

OpenStudy (anonymous):

flvs

OpenStudy (vera_ewing):

hmm ok well i can't really explain how to do it so you'll understand, but the link i gave you did a pretty good job. just look at what they said and try to understand. :)

OpenStudy (anonymous):

thx anyways though

OpenStudy (vera_ewing):

sure :)

OpenStudy (anonymous):

I'm not sure I understand what it's asking! D:

OpenStudy (anonymous):

lol um 1 sec

OpenStudy (anonymous):

Kay

OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. My Bday is 12/04/1998

OpenStudy (anonymous):

an equation like f(x)=3(x-8)2+15 but made with my bday

OpenStudy (anonymous):

Well General quadratic formula ... y = a(x-h)² + k f(x)=2(x-12)^2+4 g(x)=3(x-12)^2+4 h(x)=-1(x-12)^2+4

OpenStudy (anonymous):

ok so it would be f(x)=2(x-12)^2+4 for the first question??

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

omg ty ok so the next part is

OpenStudy (anonymous):

Using complete sentences, explain how to convert your birthday function into standard form.

OpenStudy (anonymous):

Hold on let me work on this >.< So standard form is; Standard Form is ... Ax^2 + By = C use F.O.I.L on (x - 12)^2 and combine it with +15.

OpenStudy (anonymous):

were is the 15 from??

OpenStudy (anonymous):

/).< I was talking to someone else when I wrote that lol >.< I mean 4

OpenStudy (anonymous):

oh lol ok so im foiling (x-12)^2 by 4??

OpenStudy (anonymous):

so it would be 4x-48)^2

OpenStudy (anonymous):

Nono, (x - 12)(x - 12) + 4

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

so it would be x^2-12x-12x-144+4 x^2-24x-140??

OpenStudy (anonymous):

Yep, that is your standard form equation

OpenStudy (anonymous):

ok so how would i explain that in sentences??

OpenStudy (anonymous):

hm... explain exactly what I showed you, and how you worked to get it.

OpenStudy (anonymous):

So?? Foil x minus twelve squared plus four to get x squared minus twelve x minus twelve x minus onehundred and forty four plus four. Then Simplify to x squared minus twenty four x minus one hundred and forty which is standard form>

OpenStudy (anonymous):

/).< Oh boy

OpenStudy (anonymous):

lol is that correct

OpenStudy (anonymous):

Not quite.. explain the steps, not the equations, but do include the outcome

OpenStudy (anonymous):

So more like multiply x by x ,x by 12, 12 by x, 12 by 12, which gives you x squared minus twelve x minus twelve x minus onehundred and forty four plus four. Then simplify by adding the 12 x's and subtract 4 from the 140 which gives you x squared minus twenty four x minus one hundred and forty which is standard form.

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

yes ty lol heres the next part

OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

Average rate of change is \[\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (anonymous):

ok so what would be b and what would be a??

OpenStudy (anonymous):

The equation to use is f(x) = 2(x -12)^2 + 4

OpenStudy (anonymous):

idk im still confused

OpenStudy (anonymous):

i cant figure out whats a and b

OpenStudy (anonymous):

And in this case, I'd actually do this; \[\frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\] \[y _{2} = 2(7 - 12)^{2} + 4\] \[y _{1} = 2(4 -12)^{2} + 4\]

OpenStudy (anonymous):

??where is the 7 from??

OpenStudy (anonymous):

x(1) = 4 x(2) = 7

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

Sorry, should have elaborated a bit more

OpenStudy (anonymous):

ok but what about f(b)−f(a/)b−a

OpenStudy (anonymous):

and its ok youre so helpful

OpenStudy (anonymous):

Scratch that out! Go with the second one I showed you >.<

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i subtract them

OpenStudy (anonymous):

So you have to solve the equations, before you're able to do \[\frac{ y _{2} - y _{1} }{ x _{2} -x _{1} }\]

OpenStudy (anonymous):

yeah how do i subtract y2=2(7−12)2+4 y1=2(4−12)2+4

OpenStudy (anonymous):

Solve the equations first

OpenStudy (anonymous):

wait which one??

OpenStudy (anonymous):

both of them xD

OpenStudy (anonymous):

im simplifying y2 and y1 then subtracting them right

OpenStudy (anonymous):

Yes! Simplify the equations (there are no variables no so it'll be easy to come up with a straight-forward answer)

OpenStudy (anonymous):

so subtracted it would be 12

OpenStudy (anonymous):

Okay, then what is 7-4?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so then the final answer is 4??

OpenStudy (anonymous):

12/3

OpenStudy (anonymous):

Yep :D

OpenStudy (anonymous):

omg yes youre perfect lol now how do i explain that in sentences

OpenStudy (anonymous):

How did I help you find the answer? What steps did we take to get the answer?

OpenStudy (anonymous):

lol a lot of them

OpenStudy (anonymous):

Explain the steps >.<

OpenStudy (the_fizicx99):

Hmm, Well to be honest, you can use \(\ \dfrac{f(b) - f(a)}{b - a} \) it's just a bit more tedious than change in y over change in x

OpenStudy (anonymous):

thx fiz lol

OpenStudy (anonymous):

Shhh Fizi I know we caaan but I like doing it the other way dweeb

OpenStudy (the_fizicx99):

;-; Bu-but D;

OpenStudy (anonymous):

so you subtract y2 from y1 whixh is y2=2(7−12)2+4 y1=2(4−12)2+4 to get -16-(-28)=12 then subtract 7-4=3 then divide the 12/3=4 which is the answer

OpenStudy (anonymous):

Sure, that's a very simple way to put all of those steps.. xD

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

i also have to graph it know where i can do that??

OpenStudy (anonymous):

ohh shiz I don't know.. um.. @tHe_FiZiCx99

OpenStudy (the_fizicx99):

How to graph vertex form? Pretty simple :3 I'm still reading the whole problem

OpenStudy (anonymous):

actually i have to graph the answer to number 2

OpenStudy (anonymous):

I'm not good with graphing things. it takes me like an hour to graph something -_-

OpenStudy (anonymous):

lol thats ok

OpenStudy (the_fizicx99):

Ohhh wait, it says to graph your function in step 3 . . . I think, graphing standard form is actually easier than vertex form

OpenStudy (anonymous):

lol yeah i have to graph x^2-24x-140

OpenStudy (anonymous):

just use an online graphing calculator xD

OpenStudy (anonymous):

lol i used wolphram alpha